OpenStudy (anonymous):

how do you find the the second dy/dx of x^2+y^2 = 4

OpenStudy (anonymous):

derivative wrt x right? did you find \[y'=-\frac{x}{y}\]?

myininaya (myininaya):

\[2x+2yy'=0 => x+yy'=0 =>yy'=-x= > y'=\frac{-x}{y}\]

OpenStudy (anonymous):

because that is the first step if you want \[y''\] if not let me know and we can find it quickly. but once you have that, it is the quotient rule

OpenStudy (anonymous):

ok i really meant quickly clearly!

OpenStudy (anonymous):

\[\implies\]

myininaya (myininaya):

\[y''=\frac{-1y-(-x)y'}{y^2}=\frac{-y+xy'}{y^2}=\frac{-y+x \cdot \frac{-x}{y}}{y^2}\]

OpenStudy (anonymous):

\implies

OpenStudy (anonymous):

\[=\frac{-y^2-x^2}{y^3}\]

myininaya (myininaya):

\[=\frac{-y+\frac{-x^2}{y}}{y^2}=\frac{y}{y} \frac{-y +\frac{-x^2}{y}}{y^2} =\frac{-y^2-x^2}{y^3}\] \[=\frac{-(x^2+y^2)}{y^3}=\frac{-4}{y^3}\]

OpenStudy (anonymous):

can someone help me

OpenStudy (anonymous):

where did the -4 come from?

OpenStudy (anonymous):

don't bite!

myininaya (myininaya):

\[x^2+y^2=4 => -(x^2+y^2)=-4\]

OpenStudy (anonymous):

you bit

OpenStudy (anonymous):

\implies

OpenStudy (anonymous):

\[x^2+y^2=4\implies -(x^2+y^2)=-4\]

myininaya (myininaya):

how come my latex isn't showing up nice and pretty?

OpenStudy (anonymous):

looks good to me

OpenStudy (anonymous):

but my arrows are better

myininaya (myininaya):

shhh.. i didn't bite hard did i ?

OpenStudy (anonymous):

\[x^2+y^2=4\iff -(x^2+y^2)=-4\]

OpenStudy (anonymous):

i meant "don't bite" as in "i am baiting you"

myininaya (myininaya):

lol ok

OpenStudy (anonymous):

holdon for the first dy/dx i got -2x/2y and that is right

myininaya (myininaya):

yes you cancel the two's though

OpenStudy (anonymous):

ohh lol alright i see ty